1.      (i) Use normal approximation to a Binomial Distribution (X) having n=500 trials and p = 0.025 for finding out:

(a)   P(X>10);                              (b) P(X<18)       (c) P(X>21)               (d) P(9<X<14)    (4)

        



1
Expert's answer
2021-07-05T04:12:41-0400

Binomial distribution bin(n,p) can be estimated using a normal distribution norm(np,np(1-p). However, correction for continuity should be carried out. Below is the correction factor table.

"If \,P(X=y) \,use \,P(y-9.5<X<y+0.5)"

"If \,P(X>y) \,use \,P(X>y+0.5)"

"If \,P(X<) \,use \,P(X<y-0.5)"

"If\, P(X\le) \,use\, P(X< y+0.5)"

"If \, P(X\ge) \,use \,P(X>y-0.5)"

Before using normal approximation for binomial distribution the conditions "np\ge5" or "n(1-p)\ge5" should be met.

"n=500"

"P=0.025"


"np=500\times0.025=12.5"

"n(1-p)=500\times0.9725=487.5"

Both np and n(1-p) are greater that 5. Thus, we can use normal approximation.

"np=125"

"\sqrt{np(1-p)}=3.49106"

(a) "P(X>10)"

Correcting for continuity, "P(X>10)" becomes "P(X>10.5)" based on the above table.

"P(>10.5)=1-P(X<10.5)"

"=1-P(z<\frac{10.5-12.5}{3.49106})"

"=1-P(z<-0.573)"

"=1-0.28336"

"=0.71664"

(b). "P(X<18)"

Correcting for continuity becomes "P(X<17.5)"

"P(X<17.5)=P(z<\frac{17.5-12.5}{3.49106})"

"=P(z<1.432)"

"=0.924"

(c) "P(X>21)"

Correcting for continuity, "P(X>21)" becomes "P(X>21.5)"

"P(>21.5)=1-P(X<21.5)"

"=1-P(z<\frac{21.5-12.5}{3.49106})"

"=1-P(z<2.578)"

"=1-0.995"

"0.005"

(d) "P(9<X<14)"

"P(9<X<14)=P(X<14)-P(X<9)"

Correcting for continuity, "P(9<X<14)" becomes "P(X<13.5)-P(X<8.5)"

"P(X<13.5)-P(X<8.5)=P(z<\frac{13.5-12.5}{3.49106})-P(z<\frac{8.5-12.5}{3.49106})"

"=P(z<0.286)-P(z<-1.146)"

"=0.6127-0.1259"

"=0.4868"